Circular No. 6739

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOTICE TO CONTRIBUTORS AND SUBSCRIBERS Some respondents have expressed worry that the procedure given on

IAUC 6737 for reporting CCD and photographic discoveries of supernovae and novae will incur delays in announcement that could jeopardize significant scientific results. For some years we have been routinely relaying all remotely promising reports privately to spectroscopists for confirmation, and almost every observer in the world able and willing to make these confirmations is on our list. The delay is in securing spectroscopy. By accepting discoveries on the basis of compliance with the first three requirements (accurate position, second-night confirmation, nonappearance on appropriate earlier images), we hope to be able significantly to reduce the delay—albeit at the expense of occasionally announcing observations of some unknown long-period variable star instead. With a record 109 genuine supernovae already recognized in 1997 (in addition to numerous false alarms), some organization of the activity is clearly essential if it is to remain manageable.

Proponents of the use of e-mail exploders and the WWW for the widespread dissemination of incomplete and uncertain reports suggest that such use provides more of a guarantee that “science” can be done, conducted in a timely manner, and that nothing will be lost. Even if an occasional imperfect report could be salvaged in this way, the whole process is extraordinarily inefficient and carried out with an enormous “spinning of wheels” by hundreds of participants. A lofty scientific ideal is all very well, but complaints would surely be rampant if the confusion led to misplaced credit for a discovery. Several serious observers have also mentioned that they would be embarrassed to publicize in the internet claims that might be erroneous. It has been proposed that those who disseminate discovery reports in the internet should be sure also to inform the Central Bureau; it would also help if those who do contact us were to let us know if they have also distributed their reports widely to others.

One point on IAUC 6737 referred to the unreliability of some sources of positional information about minor planets. The most longstanding electronic source for listing the minor planets in a particular region of the sky at a given time is the CBAT/MPC Computer Service, which has reliably provided this information for the numbered minor planets for well over a decade. A new version of this feature, freely available in the WWW at

http://cfa-www.harvard.edu/ps/CheckSN, also includes many unnum-

bered minor planets (although for these the data are not fully guaranteed). While a general-purpose facility, it is particularly suited for checking candidate supernovae, for it will show the minor planets in the vicinity of a specified M, NGC, IC, UGC or MCG galaxy at the time in question.

1997 September 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden